OpenStudy (anonymous):

can somebody help me with th slope of each line and then detemine whether th two lines are parrelel, perpendicular, or neither -5x+3y=3 and 2x+7y=0

OpenStudy (anonymous):

neither

OpenStudy (whpalmer4):

To get the slope, rearrange the equation in slope intercept form: \(y = mx+b\) where \(m\) is the slope. Parallel lines have = slope. Perpendicular lines have slopes such that the product = -1

OpenStudy (anonymous):

but what would be the slopes i came up with 1/3 and 2x/7 did imiss somethn

OpenStudy (whpalmer4):

\[-5x+3y=3\]add 5x to both sides\[3y=5x+3\]divide by 3\[y=\frac{5}{3}x+1\]\[m=\frac{5}3\] \[2x+7y=0\]subtract 2x from both sides\[7y=-2x\]divide by 7\[y=-\frac{2}{7}x\]\[m=-\frac{2}{7}\]

OpenStudy (whpalmer4):

the slopes aren't equal, and their product does not equal -1, so they are neither parallel nor perpendicular.

OpenStudy (anonymous):

thats whats up thank u ima be back to ask you a ? lol

OpenStudy (anonymous):

were can i get grah help from

OpenStudy (whpalmer4):

My magic decoder ring suggests that you meant to ask: "Where can I get help with graphs?" If that is true, could you be a bit more specific? Do you need help reading them, making them, or what?

OpenStudy (anonymous):

i got a problem write the qation for the line the points on here is at the(x 2 y 5)nd 0,6the equation of the line is y= type y+x+b

OpenStudy (whpalmer4):

How about you try writing that again?

OpenStudy (anonymous):

point 2,5 and 0,6 write the equation for the line use y=mx+b

OpenStudy (whpalmer4):

Okay, the equation for the slope given two points \((x_1,y_1),(x_2,y_2)\) is \[m=\frac{y_2-y_1}{x_2-x_1}\] Once you have the slope, the equation for a line with known slope \(m\) going through a point \((x_0,y_0)\) is \[y-y_0=m(x-x_0)\] and then a bit of algebra to rearrange it into the requested slope-intercept form.

OpenStudy (anonymous):

the slope would be 1/-3 rise over run up one to the left 3

OpenStudy (anonymous):

1/-2 my bad

OpenStudy (whpalmer4):

Okay, your latter value for the slope is correct.

OpenStudy (anonymous):

how would i make the problem 6-5=1 and 0-2=2 h ld i turn this into y= answer

OpenStudy (anonymous):

how would i turn this into an equation solving y=

OpenStudy (whpalmer4):

You know the slope, and you know a point on the line. Use the second formula in my post to construct the equation.

OpenStudy (anonymous):

correct me im wrong here y=6-5=1/2(0-2)

OpenStudy (whpalmer4):

You only use 1 point, not 2.

OpenStudy (anonymous):

oh so y=6x-5

OpenStudy (anonymous):

im not getting how you come up with the solution for what i have figued out

OpenStudy (whpalmer4):

Okay, you got the slope, m = -1/2, right? Now use the formula for a line given the slope and a known point: \[y-y_0 = m(x-x_0)\]Our known point is (2,5): \[y-5=-\frac{1}{2}(x-2)\]You could also do it with the other known point (0,6): \[y-6=-\frac{1}{2}(x-0)\]Now add 5 or 6 respectively to both sides and simplify.

OpenStudy (anonymous):

final answer y=-1/2+4

OpenStudy (anonymous):

y-5=2-xover 2

OpenStudy (whpalmer4):

No...your first answer has no \(x\) in it. Your second answer isn't in the proper form. \[y-5=-\frac{1}{2}(x-2)\]Add 5 to both sides. Now use the distributive property to expand the right hand side and simplify. Don't forget that -ive * -ive = +ive.